# Mil2: Efficient Cloth Simulation Using Non-distance Barriers and Subspace Reuse

### v3

@@ 5.1. Pre-computed subspace update
The most expensive computation however is the subspace projection of $\Delta\mathsf{H}$ i.e., evaluating $\mathsf{U}^{\top}\Delta\mathsf{H}\mathsf{U}$. The complexity is $O(N^{2}r)$ on the surface. This is prohibitive even on the GPU if we want the simulation to be interactive and high-resolution at the same time.
We obviate this difficulty exploiting unique structures of $\mathsf{U}^{\top}\Delta\mathsf{H}\mathsf{U}$. First, we set $\mathsf{A}_{i}=\mathsf{B}_{i}=\mathsf{Id}$ in Eq. (3) as an identity matrix ($\mathsf{Id}$) for local projection of collision constraints. Intuitively, doing so is equivalent to deploying a nonlinear spring pulling a penetrating vertex to its target (collision-free) position. With EXP, the strength of the spring grows exponentially after each LG iteration if the vertex keeps penetrating. It is immediately noticed that $\Delta\mathsf{H}$ becomes a diagonal matrix under this treatment.
@@
(19) $\mathsf{U}^{\top}\Delta\mathbf{H}\mathbf{U}=\sum_{j=1}^{k}w_{j}U_{c_{j}}\otimes U_{c_{j}}\in\mathbb{R}^{r\times r}.$
That said the change of the subspace matrix caused by those $k$ colliding vertices is the weighted summation over $k$ rank-one matrices. All of these $U_{c_{j}}\otimes U_{c_{j}}$ can be pre-computed, and every $U_{c_{j}}\otimes U_{c_{j}}$ is a $r\times r$ symmetric matrix. In practice, we only save its upper triangle at each vertex, and the corresponding memory footprint is lightweight. The computation of the summation, on the other hand, is unfolded at each of $r(r-1)/2$ matrix elements using CUDA. The pre-computed subspace update using Eq. (19) is $50$ to $150$ times faster than directly evaluating $\mathsf{U}^{\top}\Delta\mathsf{H}\mathsf{U}$ using cuBlas.
After $q$ is computed, we use A-Jacobi to relax the remaining high-frequency residual as in Eq. (17) and move to the next LG iteration till the stopping criterion is satisfied. After an LG iteration, partial CCD ensues, and $w_{i}$ of the collision constraints are updated per Eq. (7). Nevertheless, our subspace reuse scheme can efficiently calculate the updated (reduced) global matrix and allows Mil2 to enjoy the advantages of both subspace solvers and iterative solvers with minimum computational costs. Such combined efficiency and convergence are unseen in previous GPU algorithms (e.g., see Fig. 6). It is noteworthy that this method is also highly effective for volumetric deformable models, where low-frequency motions are particularly challenging for iterative solvers.
@@ 6. Residual Forwarding
In interactive applications, simulation modules normally have prescribed time budgets regardless if the solver reaches the convergence. This hard constraint forces the simulator to enter the final CCD-based linear search filtering and truncate the position update $\Delta x$ by the global TOI: $\Delta x\leftarrow\alpha t_{\text{TOI}}\cdot\Delta x$. If the TOI is a smaller quantity, severe damping or locking artifacts could be produced as a significant portion of the system energy dissipates by the early termination.
To partially alleviate this issue, we propose a post-step treatment namely residual forwarding, or RF in short. The idea of RF is to estimate the remaining residual generated by small-step line search filtering and/or non-convergent LG iterations. Recall that each time step, the simulation seeks a minimizer $x^{\star}$ of Eq. (5), which ideally should possess a vanished gradient $\nabla E(x^{\star})=0$. By the end of a time step, if LG iterations fail to fully converge, the resulting $x^{*}$ is different from $x^{\star}$ such that $x^{\star}=x^{*}+\delta x$. The gradient of the variational energy $\nabla E(x^{*})=f_{r}\neq 0$ represents unbalanced residual forces in the system. RF seeks a kind of virtual force $\delta f$ at the next time step to mitigate the damping artifacts induced by $\delta x$. That said, Eq. (2) becomes:
(20) $z=x^{*}+h\dot{x}^{*}+h^{2}\mathsf{M}^{-1}(f_{\text{ext}}+\delta f).$
@@
(21) $z=x^{\star}+h\dot{x}^{\star}+h^{2}\mathsf{M}^{-1}f_{\text{ext}}=x^{*}+h\dot{x}^{*}+2\delta x+h^{2}\mathsf{M}^{-1}f_{\text{ext}}.$
@@
(23) $\delta x=-\left(\nabla^{2}E(x^{*})\right)^{-1}\left(f_{r}+\epsilon(\|\delta x\|^{2})\right).$
@@
Taking a closer look, it is noted that computing $\left(\nabla^{2}E(x^{*})\right)^{-1}f_{r}$ is equivalent to taking one more Newton solve by the end of the previous time step, which can be efficiently approximated with subspace reuse. In RF, we do not need to perform CCD-based line search filtering, and the weights for all the collision barriers is $k$ i.e., $a$ is set zero in Eq. (7).
@@ Discussion
RF is a heuristic treatment when the current time step must end for other time-critical tasks. If $\epsilon(\|\delta x\|^{2})$ is big, RF becomes erroneous and generates severe artifacts. Honestly, there is nothing much we can do if the application aggressively restrains the available time budget for simulation since Eq. (5) is highly nonlinear by itself. Conceptually, RF moves some computation e.g., solve for $\left(\nabla^{2}E(x^{*})\right)^{-1}f_{r}$ to the next time step. The advantage of RF is skipping the line search filtering since $z$ could embody an overlapping and penetrating configuration. The collision constraint set is then updated in the follow-up LG iteration in the future time step. RF, on the other hand, allows non-colliding vertices to move under cloth momentum and elasticity even with a small $t_{\text{TOI}}$ (i.e., from the previous time step). As a result, damping/locking artifacts are ameliorated.
Figure 7. Algorithm overview.  Mil2 is an efficient cloth simulation pipeline. It leverages a subspace reuse technique to substantially improve the convergence of GPU-based iterative solvers. The costly CCD is also replaced with a non-distance-based barrier formulation i.e., EXP. The pipeline also features a residual forwarding mechanism, which generates plausible animation even with small exiting TOI. We use a bigger subspace of $\bar{r}$ dimension to obtain a good warm start at the beginning of each time step. A more compact $r$-dimension subspace is used for handling EXP-in-the-loop optimization. Each time step consists of several blocks of LG iterations, which are visualized with curved double arrows.
$z\leftarrow x^{*}+h\dot{x}^{*}+h^{2}(\mathsf{M}^{-1}f_{\text{ext}}+\delta f)$ // $\delta f$ is from RF
@@ 1 $x\leftarrow z$
3 while $\|\Delta x\|<\epsilon_{\text{initial}}$ do
@@ 7 $x^{+}\leftarrow x+\Delta x$
$\langle B,t_{\text{TOI}}\rangle\leftarrow$ FullCCD$(x^{-},x^{+})$// $B$ is the latest barrier
8 $x\leftarrow x+t_{\text{TOI}}\cdot\Delta x$
@@ 13             $B_{i}\leftarrow$PartialCCD$(x^{-},x^{+})$
      $\langle B,t_{\text{TOI}}\rangle\leftarrow$ FullCCD$(x^{-},x^{+})$ // update $B$ at each outer loop
@@
$x\leftarrow x+t_{\text{TOI}}\cdot\Delta x$// exiting line search filtering
17 if $t_{\text{TOI}}<\epsilon_{\text{TOI}}$ then
@@ 22      $\delta f\leftarrow\frac{2\delta x\mathsf{M}}{h^{2}}$
ALGORITHM 1 Mil2 simulation pipeline.
@@ 7. Simulation Pipeline
We now have all the pieces to assemble our simulator. Fig. 7 visualizes major steps along our pipeline, and the pseudocode is also outlined in Alg. 1. The pre-computation stage performs the eigendecomposition of the rest-shape global matrix $\mathsf{H}$, which is agnostic on the collision or barrier constraints. The eigendecomposition computes $\bar{r}$ eigenvectors corresponding to the smallest eigenvalues, and we typically set $\bar{r}$ as a few hundred i.e., $\bar{r}>r$. Because the rest-shape modal global matrix $\Lambda=\mathsf{U}^{\top}\mathsf{H}\mathsf{U}$ is diagonal, and the subspace projection is efficient on GPU, increasing the dimensionality of $\mathsf{U}$ from $r$ to $\bar{r}$ does not induce extra computational costs but makes the system solve (w.o. collision) more effective. The first $r<\bar{r}$ eigenvectors form the set of basis vectors $\mathsf{V}$ of a more compact subspace, and we set $r=30$ in our implementation. $\mathsf{V}$ is for subspace reuse when collision constraints are taken into account. For each vertex $j$, we pre-compute the corresponding $\mathsf{V}_{j}\otimes\mathsf{V}_{j}$ as in Eq. (19) for fast computation of the updated subspace matrix. Lastly, $\mathsf{U}^{\top}\mathsf{H}X$ and $\mathsf{V}^{\top}\mathsf{H}X$ are also pre-computed for faster assembly of the r.h.s.of the global solve (i.e., Eqs. (16) and (18)).
@@
Figure 8. Cloth twisting with EXP.  We rotate the table cloth at both ends for $2,880$ degrees. The simulation includes over $120$K DOFs. Mil2 easily completes this test at 12 FPS. Compared with distance-based barrier i.e., (Lan et al., 2022b), EXP saves about $30\%$ LG iterations (as plotted in Fig. 10).
@@
Keeping $\mathsf{V}$ condensed is helpful for both efficiency and efficacy of the simulation. From the efficiency point of view, we know $\Lambda+\mathsf{V}^{\top}\Delta\mathsf{H}\mathsf{V}$ is not diagonal. Solving a dense linear system is only feasible when the system is of low dimension. From the efficacy point of view, the subspace reuse is effective in the low-frequency realm. Barrier constraints induced by collisions and contacts can drastically change the higher-frequency landscape of the spectrum. As a result, expanding $\mathsf{V}$ to the high-frequency spectrum is not beneficial. Since we will need to save $\mathsf{V}_{i}\otimes\mathsf{V}_{i}$ for each vertex, a more compact $\mathsf{V}_{i}$ is also more memory-friendly.
@@
Each time step begins with a warm-start computation. Typically, $z$ (as in Eq. (2)) is used as the initial guess of $x$ at the current time step and for setting up the initial barrier constraints $B$. As the collision is ignored for the warm start, $\bar{r}$-dimension subspace provides a better initialization by solving $\arg\min_{x}I+\Psi$ (line 5 in Alg. 1). The decoupled computations in the modal space of span($\mathsf{U}$) make this procedure highly efficient. For instance, the warm start only takes three to five LG iterations and less than five milliseconds for a $300$K-DOF simulation. The resulting $x$ is forwarded to the full CCD procedure. The constraint list is then built, and $B(x)$ is initialized (line 9).
@@
Afterwards, the solver works on the optimization of $\arg\min_{x}I+\Psi+B$. Similar to existing barrier-in-the-loop PD algorithms (Lan et al., 2022b), we employ a two-level iteration scheme. The outer loop consists of multiple LG iterations, and the outer convergence check is based on the change of $\|x\|$ between two consecutive outer loops. That said if $\|x\|$ nearly stays the same for two outer loops, we consider the time step is completed. Each inner loop begins with a standard LG iteration. The local projections are performed for both elasticity constraints ($\Psi$) and collision constraints ($B$). At the global solve, we re-use the subspace $\mathsf{V}$ constructed from $\mathsf{H}$ to smooth high-frequency errors and pass the residual to A-Jacobi iterations. The convergence of the inner loop is also based on the variation of $\|x\|$ from the previous inner loop. After the LG is done, the partial CCD routine is invoked (line 16) to adjust the weights of collision/contact constraints. As discussed in Sec. 4, this computation boils down to computing the inner products of 3-vectors.
@@
By the end of each time step, a full CCD and an exiting line search filtering are performed. They offer the algorithmic guarantee that the simulation is free of inter-penetration. If the exiting $t_{\text{TOI}}$ is too small suggesting possible locking and overdamping due to insufficient outer loops, we trigger the residual forward, which estimates an optimal correction force $\delta f$ for the next step.
@@ 8. Experimental Results
We implemented the Mil2 pipeline on a desktop computer with an intel i7-12700 CPU and an nVidia 3090 RTX GPU. We used Spectra library for computing the eigendecomposition of the global stage matrix $\mathsf{H}$. It should be noted that Mil2 is also friendly with other parallel computing platforms – one can easily parallelize local projections using multi-threading, and multicore CPUs are well suited for the Gauss-Seidel method. Nevertheless, we only report the performance of Mil2 on the GPU.

### v4

@@ 5.1. Pre-computed subspace update
The most expensive computation is the subspace projection of $\Delta\mathsf{H}$ i.e., evaluating $\mathsf{U}^{\top}\Delta\mathsf{H}\mathsf{U}$. The complexity is $O(N^{2}r)$ on the surface. This is prohibitive even on the GPU if we want the simulation to be interactive and high-resolution at the same time.
We obviate this difficulty exploiting the unique structure of $\mathsf{U}^{\top}\Delta\mathsf{H}\mathsf{U}$. First, we set $\mathsf{A}_{i}=\mathsf{B}_{i}=\mathsf{Id}$ in Eq. (3) as an identity matrix ($\mathsf{Id}$) for local projection of collision constraints whose target (collision-free) positions are computed as in (Lan et al., 2022b). With NDB, the target positions can be achieved effectively within just a few LG iterations. We note that $\Delta\mathsf{H}$ becomes a diagonal matrix under this treatment.
@@
(19) $\mathsf{U}^{\top}\Delta\mathsf{H}\mathbf{U}=\sum_{j=1}^{k}w_{j}U_{c_{j}}\otimes U_{c_{j}}\in\mathbb{R}^{r\times r}.$
Therefore the change of the subspace matrix caused by those $k$ colliding vertices is the weighted summation over $k$ rank-one matrices. All of these $U_{c_{j}}\otimes U_{c_{j}}$ can be pre-computed, and every $U_{c_{j}}\otimes U_{c_{j}}$ is a $r\times r$ symmetric matrix. In practice, we only save its upper triangle at each vertex, and the corresponding memory footprint is lightweight. The computation of the summation, on the other hand, is unfolded at each of $r(r-1)/2$ matrix elements using CUDA. The pre-computed subspace update using Eq. (19) is $50$ to $150$ times faster than directly evaluating $\mathsf{U}^{\top}\Delta\mathsf{H}\mathsf{U}$ using cuBlas.
After $q$ is computed, we use A-Jacobi to relax the remaining high-frequency residual as in Eq. (17) and move to the next LG iteration until the stopping criterion is satisfied. After an LG iteration, partial CCD ensues, and $w_{i}$ of the collision constraints are updated per Eq. (7). Our subspace reuse scheme can efficiently calculate the updated (reduced) global matrix and allows the simulation to enjoy the advantages of both subspace solvers and iterative solvers with small computational costs. Such combined efficiency and convergence are unseen in previous GPU algorithms (e.g., see Fig. 6). In practice, we use two subspaces to handle simulation without and with collisions (as detailed in Sec. 8.1). It is noteworthy that this method is also highly effective for volumetric deformable models, where low-frequency motions are particularly challenging for iterative solvers.
@@ 6. Residual Forwarding
In interactive applications, simulation modules normally have prescribed time budgets regardless if the solver reaches the convergence. This hard constraint forces the simulator to enter the final CCD-based linear search filtering and truncate the position update $\Delta x$ by the global TOI: $\Delta x\leftarrow\alpha t_{TOI}\cdot\Delta x$. If the TOI is a smaller quantity, severe damping or locking artifacts could be produced as a considerable portion of the system energy dissipates by the early termination.
To partially alleviate this issue, we propose a post-step treatment namely residual forwarding, or RF in short. The idea of RF is to estimate the remaining residual generated by small-step line search filtering and/or non-convergent LG iterations. Recall that each time step, the simulation seeks a minimizer $x^{\star}$ of Eq. (5), which ideally should possess a vanished gradient $\nabla E(x^{\star})=0$. By the end of a time step, if LG iterations fail to fully converge, the resulting $x^{*}$ is different from $x^{\star}$ such that $x^{\star}=x^{*}+\delta x$. The gradient of the variational energy $\nabla E(x^{*})=-f_{r}\neq 0$ represents unbalanced residual forces in the system. RF seeks the virtual force $\delta f$ at the next time step to mitigate the damping artifacts induced by $\delta x$. Therefore, Eq. (2) becomes:
(20) $z=x^{*}+h\dot{x}^{*}+h^{2}\mathsf{M}^{-1}(f_{{ext}}+\delta f).$
@@
(21) $z=x^{\star}+h\dot{x}^{\star}+h^{2}\mathsf{M}^{-1}f_{{ext}}=x^{*}+h\dot{x}^{*}+2\delta x+h^{2}\mathsf{M}^{-1}f_{{ext}}.$
@@
(23) $\delta x=\left(\nabla^{2}E(x^{*})\right)^{-1}\left(f_{r}-\epsilon(\|\delta x\|^{2})\right).$
@@
Taking a closer look, it is noted that computing $\left(\nabla^{2}E(x^{*})\right)^{-1}f_{r}$ is equivalent to taking one more Newton solve by the end of the previous time step, which can be efficiently approximated with subspace reuse. In RF, we do not need to perform CCD-based line search filtering, and the weights for all the collision constraints is $k$ i.e., $a_{i}$ for $i$-th collision constraint is set zero in Eq. (7).
@@ Discussion
RF is a heuristic treatment when the current time step must end for other time-critical tasks. If $\epsilon(\|\delta x\|^{2})$ is big, RF becomes erroneous and generates artifacts. Honestly, there is nothing much we can do if the available time budget is aggressively restrained. Conceptually, RF moves some computation e.g., solve for $\left(\nabla^{2}E(x^{*})\right)^{-1}f_{r}$ to the next time step. The advantage of RF is skipping the line search filtering since $z$ could embody an overlapping and penetrating configuration. The collision constraint set is then updated in the follow-up LG iteration in the future time step. RF, on the other hand, allows non-colliding vertices to move under cloth momentum and elasticity even with a small $t_{TOI}$ (i.e., from the previous time step). As a result, damping/locking artifacts are ameliorated.
Figure 7. Algorithm overview.  Our method leverages a subspace reuse technique to improve the convergence of GPU-based iterative solvers. A large portion of costly CCDs is replaced with a non-distance non-distance barrier formulation i.e., NDB. The pipeline also features a residual forwarding mechanism, which generates plausible animation even with small exiting TOI. We use a bigger subspace of $\bar{r}$ dimension to obtain a good warm start at the beginning of each time step. A more compact $r$-dimension subspace is used for handling NDB-in-the-loop optimization. Each time step consists of several blocks of LG iterations, which are visualized with curved double arrows.
$z\leftarrow x^{*}+h\dot{x}^{*}+h^{2}(\mathsf{M}^{-1}f_{ext}+\delta f)$ // $\delta f$ is from RF
@@ 1 $x\leftarrow z$
3 while $\|\Delta x\|<\epsilon_{initial}$ do
@@ 7 $x^{+}\leftarrow x+\Delta x$
$\langle B,t_{TOI}\rangle\leftarrow$ FullCCD$(x^{-},x^{+})$// $B$ is the latest barrier
8 $x\leftarrow x+t_{TOI}\cdot\Delta x$
@@ 13             $B_{i}\leftarrow$PartialCCD$(x^{-},x^{+})$
      $\langle B,t_{TOI}\rangle\leftarrow$ FullCCD$(x^{-},x^{+})$ // update $B$ at each outer loop
@@
$x\leftarrow x+t_{TOI}\cdot\Delta x$// exiting line search filtering
17 if $t_{TOI}<\epsilon_{TOI}$ then
@@ 22      $\delta f\leftarrow\frac{2\delta x\mathsf{M}}{h^{2}}$
ALGORITHM 1 Our simulation pipeline.
@@ 7. Simulation Pipeline
We now have all the pieces to assemble our simulator. Fig. 7 visualizes major steps along our pipeline, and the pseudocode is also outlined in Alg. 1. The pre-computation stage performs the eigendecomposition of the rest-shape global matrix $\mathsf{H}$, which ignores the collision constraints. We use two subspaces to handle global solves without and with collisions. Specifically, a subspace of higher dimension is first constructed out of $\bar{r}$ eigenvectors. Because the rest-shape modal global matrix $\Lambda=\mathsf{U}^{\top}\mathsf{H}\mathsf{U}$ is diagonal, and the subspace projection of r.h.s. vector is efficient on GPU, increasing the dimensionality of the subspace for collision-free global solve is worthy and effective. The first $r<\bar{r}$ eigenvectors of $\mathsf{U}$ form a smaller set of bases $\mathsf{V}$ of a more compact subspace, and we set $r=30$ in our implementation. $\mathsf{V}$ is for subspace reuse when collision constraints are taken into account. For each vertex $j$, we pre-compute the corresponding ${V}_{j}\otimes{V}_{j}$ as in Eq. (19) for fast computation of the updated subspace matrix. Lastly, $\mathsf{U}^{\top}\mathsf{H}X$ and $\mathsf{V}^{\top}\mathsf{H}X$ are also pre-computed for faster assembly of the r.h.s. of the global solve (i.e., Eqs. (16) and (18)).
@@
Figure 8. Cloth twisting NDB.  We rotate the table cloth at both ends for $2,880$ degrees. The simulation includes over $120$K DOFs. Our method completes this test at 12 FPS. Compared with the distance-based barrier weighted PD method i.e., (Lan et al., 2022b). NDB saves about $30\%$ LG iterations (as plotted in Fig. 9).
@@
Keeping $\mathsf{V}$ condensed is helpful for both efficiency and efficacy of the simulation. From the efficiency point of view, we know $\Lambda+\mathsf{V}^{\top}\Delta\mathsf{H}\mathsf{V}$ is not diagonal. Solving a dense linear system is only feasible when the system is of low dimension. From the efficacy point of view, the subspace reuse is effective in the low-frequency realm. Barrier constraints induced by collisions and contacts can drastically change the higher-frequency landscape. As a result, expanding $\mathsf{V}$ to the high-frequency spectrum is not beneficial. Since we will need to save ${V}_{j}\otimes{V}_{j}$ for each vertex, a more compact ${V}_{j}$ is also more memory-friendly.
@@
Each time step begins with a warm-start computation. Typically, $z$ (i.e., in Eq. (20)) is used as the initial guess of $x$ at the current time step and for setting up the initial constraints $B^{NDB}$. As the collision is ignored for the warm start, $\bar{r}$-dimension subspace provides a better initialization by solving $\arg\min_{x}I+\Psi$ (line 5 in Alg. 1). The decoupled computations in the modal space of span($\mathsf{U}$) make this procedure highly efficient. For instance, the warm start only takes three to five LG iterations and less than five milliseconds for a $300$K-DOF simulation. The resulting $x$ is forwarded to the full CCD procedure. The constraint list is then built, and $B(x)$ is initialized (line 9).
@@
Afterwards, the system solves for the optimization of $\arg\min_{x}I+\Psi+B$. Similar to existing barrier-in-the-loop PD algorithms (Lan et al., 2022b), we employ a two-level iteration scheme. The outer loop consists of multiple LG iterations, and the outer convergence check is based on the norm of the change of $x$ between two consecutive outer loops. If we uniformly scale the cloth meshes to a unit size, $\|\Delta x\|\leq 1E-3$ is a good choice for convergence check. For simulations involving fast-moving objects, setting $\|\Delta x\|\leq 5E-4$ may be needed. Each inner loop begins with a standard LG iteration. The local projections are performed for both elasticity constraints ($\Psi$) and collision/barrier constraints ($B$). At the global solve, we re-use the subspace $\mathsf{V}$ to smooth high-frequency errors and pass the residual to A-Jacobi iterations. The convergence of the inner loop is also based on $\|\Delta x\|$ from the previous inner loop ($\|\Delta x\|\leq 5E-2$). The partial CCD routine is then invoked (line 16) to adjust the weights of collision/contact constraints. As discussed in Sec. 4, this computation boils down to computing the inner products of 3-vectors.
@@
By the end of each time step, a full CCD and an exiting line search filtering are performed. They offer the algorithmic guarantee that the simulation is free of inter-penetration. If the exiting $t_{TOI}$ is too small suggesting possible locking and overdamping due to insufficient outer loops, we trigger the residual forward, which estimates an optimal correction force $\delta f$ for the next step.
@@ 8. Experimental Results
We implemented the proposed simulation framework on a desktop computer with an intel i7-12700 CPU and an nVidia 3090 RTX GPU. We used Spectra library for computing the eigendecomposition of the global stage matrix $\mathsf{H}$. It should be noted that our method is also friendly with other parallel computing platforms – one can easily parallelize local projections using multi-threading, and multicore CPUs are well suited for the Gauss-Seidel method. Nevertheless, we only report the performance on the GPU. The reader can find animated results in the accompanying video demo.